mathslover (mathslover):

Logarithm Question... Given \(n^3 < 10^n\) for a fixed positive integer \( n \ge 1\) , prove that \((n+1)^3 < 10^{n+1}\)

mathslover (mathslover):

Regarding this question, I tried : \( \log n^3 < n \) \(\log n^3 +1 < n + 1\) \((\log n^3 + 1)^3 < (n+1)^3\) \((\log(10n^3) )^3 < (n+1)^3\) Not sure, whether it is the best way to do it, but, I think, we can get something from it.

OpenStudy (anonymous):

induction hehe

mathslover (mathslover):

And, if I just don't add 1 both sides and follow other method : \(\log (n^3) < n \\ 3\log (n) < n \\ 3 < \cfrac{n}{\log n}\) Do I get something from here? Well, i'm not sure.

OpenStudy (anonymous):

why dnt u use induction :O

OpenStudy (anonymous):

or suppose to use log*

mathslover (mathslover):

Well, I would love to see how to do it with induction, but as log is more important for me (currently) , so I'm aiming it to solve by using Logarithm Properties.

mathslover (mathslover):

I got something : \(\large n < 10^{n + \cfrac{1}{3}}\)

OpenStudy (shubhamsrg):

I see n^3 < 10^n for any n. :| Proving this much should be sufficient.

mathslover (mathslover):

For any n? Hmm... nope :/

OpenStudy (vishweshshrimali5):

put, m = n+1

OpenStudy (vishweshshrimali5):

Now, for any \(n \ge 1\), \(m \ge 2\).

OpenStudy (vishweshshrimali5):

Now, \((m^3 < 10^m)\) Since, m satisfies the condition \(m \ge 1\)

OpenStudy (vishweshshrimali5):

Now, again put m = n+1

OpenStudy (vishweshshrimali5):

You will get your required proof.

mathslover (mathslover):

Oh :P .... I thought there will be a use of Logarithm... !

OpenStudy (anonymous):

for me i would try this :- for n=1 its true lets it be true for n show for n+1 (n+1)^3=n^3+3n^2+3n+1 =\( n^3(1+\frac{3}{n}+\frac{3}{n^2}+\frac{1}{n^3})\) the induction step n^3<10^n eq 1 1<\(1+\frac{3}{n}+\frac{3}{n^2}+\frac{1}{n^3}\le\) 8 <10 eq 2 so (n+1)3<10^n+1

OpenStudy (vishweshshrimali5):

No problem

Parth (parthkohli):

@vishweshshrimali5 But that is still induction...

Parth (parthkohli):

You can take the log of both sides and then choose \(m = n + 1\), but it makes on difference. So I guess your brother's technique is fine. :P

OpenStudy (vishweshshrimali5):

Tell you what, leave that method. Its wrong.

OpenStudy (vishweshshrimali5):

It never says that n is a general value. It can be some particular n.

OpenStudy (vishweshshrimali5):

Use only log there.

mathslover (mathslover):

Yeah I knew it already :P xD

OpenStudy (kainui):

Yeah I just noticed that as well @vishweshshrimali5 because if it's for any integer n>1, then n+1 is technically already given to you and there's nothing to prove lol.

OpenStudy (vishweshshrimali5):

Yeah @Kainui, that's why it sounded fishy to me. ;)

mathslover (mathslover):

Glad to know that the question is not that easy :P @vishweshshrimali5 , @Kainui - the methods I used above, would they help us in leading to the answer? or they were just waste :P

OpenStudy (kainui):

I can't tell you until I figure it out, seems helpful though lol

mathslover (mathslover):

Yep, sure, take your time.

OpenStudy (vishweshshrimali5):

Ok I think I got it.

OpenStudy (vishweshshrimali5):

From the first condition I get: \[3 \times \log{n} < n\] \[\implies 3\times \log{n} - n < 0\] ------------ (1)

OpenStudy (vishweshshrimali5):

Now, see that : \[( 1+ \cfrac{1}{n}) > 1\] As well as \[ 2 \ge (1 + \cfrac{1}{n})\] So, \[8 \ge (1 + \cfrac{1}{n})^3 > 1\]

OpenStudy (vishweshshrimali5):

Now, thus, \[(1 + \cfrac{1}{n})^3 < 10\] \[\implies 3 \times \log{(1+\cfrac{1}{n})} < 1\] \[\implies 3 \times \log{(\cfrac{n+1}{n})} < 1\] \[\implies 3 \times \log{n+1} - 3 \times \log{n} < 1\]

OpenStudy (vishweshshrimali5):

Let this be equation (2)

OpenStudy (vishweshshrimali5):

Adding equation 1 and 2

OpenStudy (vishweshshrimali5):

I get, \[3 \times \log{(n+1)} - n < 1\] \[\implies 3 \times \log{(n+1)} < n+1\] \[\implies (n+1)^3 < 10^{n+1}\] HENCE PROVED HURRAY !!!!

mathslover (mathslover):

From where did you get this : \(\left(1+\cfrac{1}{n}\right) ^3 < 10\) ?

OpenStudy (vishweshshrimali5):

Well, if \(2\ge(1+\cfrac{1}{n})\) then clearly cubing both sides I will get, \[8 \ge (1+\cfrac{1}{n})^3\] Now, if a number is equal to or less than 8, then it must be less than 10 so, I get the required expression.

OpenStudy (vishweshshrimali5):

The condition \(n \ge 1\) was very important if I wanted to make the argument that \[(1+\cfrac{1}{n}) > 1\]

OpenStudy (vishweshshrimali5):

Got that ?

mathslover (mathslover):

o.O yeah, I thought the signs were reversed :P

OpenStudy (vishweshshrimali5):

Nope

mathslover (mathslover):

Yep.. got it now. Thanks a lot @vishweshshrimali5 - Awesome work...! :-)

OpenStudy (vishweshshrimali5):

A simple but nice question. Hats off to the one who made that question. It was my pleasure @mathslover

OpenStudy (vishweshshrimali5):

Well actually, I am not wearing a hat right now. :P So, forgot that one.

OpenStudy (vishweshshrimali5):

*forget

mathslover (mathslover):

:) Thanks everyone.